OpenStudy (anonymous):

Solve the radical, and check all proposed solutions:

OpenStudy (anonymous):

\[\sqrt{18x-9}=x+4\]

hero (hero):

Square both sides first. Let me know what you get

OpenStudy (anonymous):

\[18x-9=x ^{2}+16\]

hero (hero):

You squared the right side incorrectly

hero (hero):

You were supposed to get \[18x-9=(x+4)^2\]

hero (hero):

Now, expand the right side.

OpenStudy (anonymous):

18x-9=(x+4)(x+4)?

hero (hero):

Okay, now multiply the right side.

OpenStudy (anonymous):

x^2+8x+16

OpenStudy (anonymous):

\[0=x ^{2}-10x+25\]

hero (hero):

Okay, now factor the right side.

hero (hero):

What happened?

OpenStudy (anonymous):

It froze up..0=(x-5)(x-5)

hero (hero):

Which is the same thing as \[0 = (x-5)^2\] Now take the square root of both sides.

OpenStudy (anonymous):

What?

hero (hero):

Do you know what a square root is?

hero (hero):

Take the radical of both sides.

hero (hero):

\[\sqrt{0} = \sqrt{(x-5)^2}\]

hero (hero):

Don't tell me you've never seen that before.

OpenStudy (anonymous):

Oh, Okay, I have, but it just looked wrong

hero (hero):

Okay, so what is the result of that?

OpenStudy (anonymous):

15?